The sum of the squares of two consecutive even positive integers is 340.Find the integers.
Answer by checkley77(12844)

x^2+(x+2)^2=340
x^2+x^2+4x+4=340
2x^2+4x+4-340=0
2x^2+4x-336=0
2(x^2+2x-168)=0
2(x+14)(x-12)=0
x-12=0
x=12 ans.
x+2=14 ans.
Proof:
12^2+14^2=340
144+196=340
340=340